Jack was to be paid $24 for raking the leaves in his yard.Jack raked the leaves in three fourths of the yard.Then Jill came over. Jill agreed to help Jack rake the rest of the yard.Jack and Jill each did the same amount of work on that part of the yard.For Jack and Jill to receive the same rate of pay, how much of the $24 should Jill be paid?
a.$1 b.$2 c.$3 d.$4 e.$6

Answer by mananth(16946)

let the yard have x units
he rakes 3/4 of it = 3x/4.
so balance will be x-3x/4 = x/4.
..
in the x/4 part half is done by jack and half by jill
so 1/2 of x/4 = x/8
..
Jack does 3x/4 +x/8 of the yard
=6x+x /8= 7x/8
Jill does x/8
The ratio of their jobs( jack : Jill) is 7:1
Jill should be paid 1/8th of 24 = $3